OpenStudy (anonymous):

Solve 4^(x^(2)-1) =8^-x logarithms or your calculator. The larger of the two solutions that you get is

OpenStudy (anonymous):

without using logarithms or your calculator

OpenStudy (watchmath):

Notice that \(4=2^2\) and \(8=2^3\) So the equation can be rewritten as \(2^{(2(x^2-1))}=2^{(2(-x))}\) \(2^{2x^2-2}=2^{-2x}\) Since the two have the same base, the exponents must be equal \(2x^2-2=-2x\) \(2x^2+2x-2=0\) \(2(x^2+x-1)=0\) Now solve \(x^2+x-1=0\) by using quadratic formula :).